Operations Report – February 6th

Crew 188 Operational Report 06/02/2018

Sol 9

Zac Trolley

Non-Nominal Systems: electrical system, electrical cords

Notes on non-nominal systems: See individual systems at the end of the report.

Generator (hours run): 15.2h

Solar: SOC 74%

Diesel: 75%

Propane: 61 psi

Ethanol Free Gasoline (5 Gallon containers for ATV): 3 Gallons

Water (trailer): 0 Gallons

Water (static): 170 Gallons

Trailer to Static Pump used: No

Water (loft) – Static to Loft Pump used: Yes

Meter 130825.4

Toilet tank emptied: Yes

ATVs Used: The three blue ATVs were used today

Oil Added: None

ATV Fuel Used: 2 Gallons

# Hours the ATVs were Used today: 2

Notes on ATVs: It is very difficult to see the fuel level in the ATVs with the suit helmet on. Also, the gas can that had a little fuel in the bottom imploded due to temperature change. There was no leak, but the structural integrity of the can has been compromised and it should be destroyed. I suggest a new location be found to keep the cans out of the sun.

Deimos rover used: No

Hours: Hours: 108.1

Beginning charge: N/A

Ending charge: N/A

Currently charging: No

Sojourner rover used: ASSIGNED TO DIRECTOR


Beginning charge:

Ending charge:

Currently charging:

Spirit rover used: No

Hours: 18.0h

Beginning charge: N/A

Ending charge: N/A

Currently charging: No

Opportunity rover used: Yes

Hours: 19.0h

Beginning charge: 100%

Ending charge: 97%

Currently charging: Yes

Curiosity rover used: Yes

Hours: 11.4h

Beginning charge: 100%

Ending charge: 99%

Currently charging: No

HabCar: Yes, to get a load of water.

General notes and comments:

During an EVA, a plastic handle from the passenger side roll cage on Demos came off. We have the handle, but no hardware to re-attach it.

Generator was turned off at 9:00

Generator was turned on at 18:04

I’ve found two extension cords that have been damaged and need to be replaced. If there are 3 prong plugs available, I can repair the cords. Please advise.

Summary of internet: Tests were run on the internet connectivity and passed on to Bernard.

Summary of suits and radios: Nominal

Summary of Hab operations: We are running low of vegetables, and there is a concern that our vegetarians will not have enough food for the remainder of the mission.

Summary of EVA operations: Nominal

Summary of GreenHab operations: The temperature gauge display has some non-functional LCD areas, a new one has been ordered.

Summary of Science Dome operations: The science dome is awaiting a portable heater as per the MDRS Director.

Summary of RAM operations: Not Operational.

Summary of health and safety issues: Nominal

Questions, concerns and requests to Mission Support: We could use another bag of toilet deodorizer, as we only have one puck left.

The “tunnel” had a few sharp edges of the metal at the corners, and a jacket was torn. I’ve taped over the edges so it doesn’t happen again.

Copyright © The Mars Society. All rights reserved. | Main Site